Evaluate ∫∫∫T 2xy dx dy dz where T is the solid in the first octant bounded above by the cylinder z = 4 − x^2 below by the x, y-plane, and on the sides by the planes x =0, y = 2x and y = 4. Answer: ∫ (4, 0) ∫ (y/2, 0) ∫ (4−x^2, 0) 2xy dz dx dy = ∫ (2, 0) ∫ (4, 2x) ∫ (4−x^2, 0) 2xy dz dy dx = 128/3
Answer #1

son 111 22g da dy dz first octant ay-plane - 2=0. 7=4-02 x=0; y = 2 ; g = 4 -420 за ла - be 2) - (2547 [OLTE 442] 351 say do